# Disentangling Core and Edge Mechanisms of the L-mode Density Limit in DIII-D

### v2

@@ Abstract
The L-mode density limit is re-examined on the DIII-D tokamak using a negative triangularity configuration to study the underlying physics. We find the limit may not be a singular disruptive boundary but a multifaceted density saturation phenomenon governed by distinct core and edge transport mechanisms. Sustained, non-disruptive operation is achieved at densities up to 1.8 times the Greenwald limit ($n_{\mathrm{G}}$) until the termination of auxiliary heating. Systematic power scans reveal distinct power scalings for the core ($n_{e}\propto P_{\mathrm{SOL}}^{0.27\pm 0.03}$) and edge ($n_{e}\propto P_{\mathrm{SOL}}^{0.42\pm 0.04}$) density limits. The edge density saturation is triggered abruptly by the onset of a non-disruptive, high-field side radiative instability that clamps the edge density below $n_{\mathrm{G}}$. In contrast, the core density continues to rise until it saturates at a level featured by substantially enhanced core turbulence. Core transport evolves from a diffusive to an intermittent, avalanche-like state, as indicated by heavy-tailed probability density functions (kurtosis $\approx 6$), elevated Hurst exponents, and a $1/f$-type power spectrum. These findings demonstrate that the L-mode density limit is determined by a combination of edge radiative instabilities and core turbulent transport. This distinction provides separate targets for control strategies aimed at extending the operational space of future fusion devices.
@@ I Introduction
High plasma density is a fundamental requirement for achieving a self-sustaining fusion reaction, as fusion power scales strongly with density [1]. However, tokamak operation is constrained by an operational density limit [2, 3, 4]. Exceeding this critical density leads to a rapid loss of confinement, often culminating a plasma termination. Overcoming this long-standing obstacle is therefore a critical step toward realizing economically viable fusion energy.
@@
The conventional metric for the density limit is the empirical Greenwald scaling [2, 3]: $n_{\text{G}}\,(10^{20}\,\mathrm{m^{-3}})=\frac{I_{\text{p}}\,(\text{MA})}{\pi a^{2}\,(\mathrm{m^{2}})}$, where $I_{\text{p}}$ denotes the plasma current and $a$ the minor radius. This scaling is notably independent of heating power. This paradigm is increasingly challenged by observations that the achievable density scales favorably with heating power, particularly in the low-confinement (L-mode) regime. Recent theoretical modeling [5, 6] and database analyses [7, 8] suggest that even in L-mode, the achievable density limit may scale positively with increased heating power, potentially exceeding the conventional Greenwald density. Furthermore, the power-dependent density limits have also been observed in stellarators [9, 10], which lack plasma current, suggesting a more universal physics based on power balance that transcends specific magnetic configurations. Therefore, further dedicated investigations are needed to develop a physics-based understanding of the fundamental mechanisms governing density limits in magnetically confined plasmas.
@@
The underlying physics of the L-mode density limit involves a complex interplay of several processes. Turbulent transport [11, 12, 13, 14, 15], radiative power losses [16, 7, 4], and magnetohydrodynamic (MHD) instabilities [17, 18] have all been reported as key contributors. These processes do not act in isolation; instead, they are coupled synergistically [3] and ultimately constrain the plasma density. Consequently, developing predictive models of density limits requires disentangling their individual roles and interdependencies.
@@
To elucidate the fundamental physics of the L-mode density limit, it is advantageous to use a plasma configuration that approximates L-mode edge conditions over a wide range of operational parameters. This avoids the complexities introduced by the L-H transition, edge transport barriers (ETBs), and edge localized modes (ELMs), which are characteristic of conventional H-mode plasmas. The negative triangularity (NT) plasmas [19, 20], characterized by an inverted “D”-shaped cross-section (see Fig. 1), provides an ideal platform for this study. These plasmas exhibit improved core confinement [21, 22, 23, 24] while maintaining L-mode-like edge conditions. While the NT edge is not identical to that of a conventional positive triangularity L-mode, its lack of a strong H-mode ETBs and ELMs makes it an excellent proxy for isolating the core and edge transport phenomena that govern density limits without confounding factors.
@@
This paper presents a re-examination of the underlying physics governing the L-mode density limit using NT plasmas on the DIII-D tokamak. We report the achievement of sustained, non-disruptive high-density operation, where the plasma density can surpass the conventional Greenwald density by a factor of 1.8 at substantial input power. A systematic power scan reveals disparate scaling dependencies for the core ($n_{e}\propto P^{0.27\pm 0.03}_{\text{SOL}}$) and the separatrix ($n_{e}\propto P^{0.42\pm 0.04}_{\text{SOL}}$) regions. Our detailed experimental measurements reveal that different mechanisms determine density limits in the core and edge plasmas. At the plasma edge, the density is constrained substantially by enhanced turbulence driven by radiative instabilities and edge cooling. In contrast, the core density is not directly limited by the radiative cooling but eventually saturates due to the emergence of enhanced, avalanche-like turbulence. These findings challenge the traditional paradigm of a single disruptive limit, advancing a more nuanced picture where distinct processes govern density saturation in different plasma regions.
@@ II Experimental Arrangement
The experiment was conducted on the DIII-D tokamak [25], with auxiliary heating power varied provided by neutral beam injection (NBI) and varied between discharges. All discharges employed a lower single-null configuration with an open divertor geometry, and the ion $\mathbf{B}\times\nabla\mathbf{B}$ drift was directed away from the primary divertor, as illustrated in Fig. 1. The plasma shape was characterized by negative triangularity at both the top and bottom ($\delta_{\text{top}}=-0.3$ and $\delta_{\text{bot}}=-0.62$), moderate elongation ($\kappa\approx 1.3$), and an edge safety factor of $q_{95}=4.1$. This paper focuses specifically on discharges with toroidal field $B_{t}=2.0$ T and plasma current $I_{p}=0.6$ MA. Additional experiments exploring different plasma current and toroidal field combinations will be reported separately [26]. In each discharge, plasma density was gradually increased via gas puffing from the top of the mainchamber. Consistent with the behavior of NT plasmas, no edge localized modes were observed in any discharge.
@@
A comprehensive suite of diagnostics systems monitored the evolving plasma state, including its equilibrium profiles, radiated power, and turbulent fluctuations. Figure 1 shows the locations of the key diagnostics. Electron density and temperature in the main chamber were measured using the Thomson scattering (TS) system [27]. Charge exchange recombination (CER) spectroscopy [28] measured the temperature, density, and toroidal and poloidal rotation velocities of carbon ions. These impurity ion profiles were used to compute radial electric field profiles ($E_{r}$) using the carbon ion force balance equation. Bolometer arrays [29] provided measurements of the radiated power distribution and total radiated power.
Several diagnostics were used to probe density fluctuations. A multichannel $\text{CO}_{2}$ interferometer [30] measured the line-averaged density and its low-$k$ fluctuations along a horizontal midplane chord and three vertical chords at major radii of $R=1.48\,\mathrm{m}$ (high-field side), $R=1.94\,\text{m}$ (core region), and $R=2.10\,\text{m}$ (low-field side). The radial interferometer polarimeter (RIP) [31] also provided line-integrated measurements of electron density fluctuations at the mid-plane. A 2D array of high-sensitivity beam emission spectroscopy (BES) [32] measured ion-scale ($k_{\theta}<3\,\mathrm{cm^{-1}}$) density fluctuations at the edge. Two 8-channel Doppler back-scattering (DBS) systems [33, 34] measured turbulent flows and low-to-intermediate-$k$ ($2<k_{\theta}<10\,\mathrm{cm^{-1}}$) density fluctuations. The radial locations and corresponding wavenumbers of the DBS measurements were calculated using the 3D ray-tracing code GENRAY [35].
@@ III.1.1 Density Saturation Beyond the Greenwald Limit
An important result of this study is demonstration of the sustained L-mode plasma operation at densities exceeding the conventional Greenwald density limit. This is shown in Fig. 2, which presents the time evolution of key parameters in two representative discharges with different levels of auxiliary heating from NBI. Across the experiment, the NBI power was varied from 3 MW up to 13 MW. As shown in Figs. 2(b) and 2(c), both core and separatrix densities increase with input power. The core density at $\rho\approx 0.7$, which is similar to the line-averaged density, consistently exceeds the Greenwald density ($n_{\text{G}}$). Crucially, the discharges with medium to high power NBI ($P_{\text{NBI}}>4$ MW) did not disrupt upon reaching their maximum density. Instead, they entered a phase of sustained saturated density that persisted until the termination of auxiliary heating.
@@
The radiated power in the divertor (Fig. 2(d)) and the main chamber (Fig. 2(e)) also increase with density initially, indicating enhanced radiative losses. The spikes in the divertor radiation power (Fig. 2(d)) are associated with the development of a radiation front that propagates from the divertor towards the high-field side (HFS).
The line-averaged density, inferred from Thomson scattering profiles, reached $1.8\times n_{\text{G}}$ in the high-power discharge. For comparison, the CO2 interferometer measured a line-averaged density of $2.0\times n_{\text{G}}$ [20, 26]. This discrepancy arises because the interferometer’s measurement includes the substantial scrape-off layer density that is presented during high-density operation. Therefore, Thomson scattering measurements are used for the analysis in this study.
III.1.2 Distinct Power Scaling for Core and Edge Density Limit
To explore the relationship between input power and achievable density, we examine the dependence of the maximum density on the power flowing into the scrape-off layer (SOL), $P_{\mathrm{SOL}}$. Here, $P_{\mathrm{SOL}}=P_{\text{in}}-P_{\text{rad,core}}-\frac{\mathrm{d}W}{\mathrm{dt}}$, accounting for total input power, core radiated power, and changes in stored energy. Averaged values over a 100 ms window are used.
@@ III.1.2 Distinct Power Scaling for Core and Edge Density Limit
Figure 3 shows different scaling relationships for the core ($\rho=0.7$) and at the separatrix. The density at $\rho=0.7$ is chosen to represent the core region, as it is close to the line-averaged density measured by the Thomson scattering. The maximum core density increases with a relatively weak power dependence, scaling as $n_{e,\text{core}}\propto P^{0.27\pm 0.03}_{\mathrm{SOL}}$. This is consistent with L-mode density limit scalings observed on other devices such as ASDEX Upgrade [8]. In contrast, the density near the separatrix exhibits a stronger dependence, scaling as $n_{e,\text{sep}}\propto P^{0.42\pm 0.04}_{\mathrm{SOL}}$. This power dependence is reminiscent of the power-dependent density limit observed in stellarators [9, 10], suggesting a more universal physical basis involving power balance and transport.
These disparate scalings are a key finding, as they imply that distinct physical processes govern density saturation in the core and edge. The remainder of this paper will investigate these processes, proposing that turbulence dominates the core density limit, while radiative dynamics are critical at the edge.
III.2 Temporal Evolution of Profiles Approaching the Density Limit
III.2.1 Profiles Evolution Towards Density Limit
@@
To investigate the underlying physical mechanisms, the temporal evolution of key plasma profiles is examined as each discharge approaches the density limit. Figure 4 displays the evolution of electron density ($n_{e}$), electron and ion temperature ($T_{e}$, $T_{i}$), and toroidal rotation rate ($\omega_{\phi}$) for a high-power discharge. As the overall density is ramped up (indicated by the color change from yellow to purple), several critical changes emerge.
While the density profile (Fig. 4(a)) rises across the plasma radius, it remains relatively flat in the core, indicating persistent particle transport. Concurrently, both the electron and ion temperature profiles (Figs. 4(b) and 4(c)) decrease substantially. A prominent feature is the flattening of the $T_{e}$ profile at the edge, a classic sign of intensified edge cooling. The toroidal rotation profile (Fig. 4(d)), which is initially peaked in the core, collapses as the density increases, leading to a significant reduction in core velocity shear.
@@
Figure 5 shows the evolution of the edge radial electric field profile ($E_{r}$) at the outer mid-plane in the high-power discharge, derived from the carbon ion force balance equation. The $E_{r}$ profile initially peaks in the core region ($\rho<0.8$), but collapses progressively as the density increases. The edge $E_{r}$ profile remains relatively flat throughout the density ramp-up. This weakening of $E_{r}$ is primarily attributed to the reduced core toroidal rotation discussed previously.
These profile changes point towards a degradation in plasma confinement. The collapse of the core rotation and the associated $E_{r}\times B$ velocity shear likely weakens the turbulence suppression. This loss of flow shear facilitates the enhanced turbulent transport that ultimately limits the core density, a topic explored in subsequent sections.
@@ III.2.2 Divergent Evolution of Turbulence in Different Regions
As the plasma approaches density saturation, a clear divergence in turbulent behavior emerges between different plasma regions. Figure 6 illustrates this by tracking electron density and temperature alongside multichannel turbulence measurements during a high-power discharge. The vertical dashed line at approximately 3700 ms marks the onset of this divergence, which features enhanced turbulence at the HFS and the core, likely due to radiative cooling. This critical event triggers distinct turbulence responses.
Initially, the local density increases across the radius (Fig. 6(a)). Concurrently, the local intermediate-$k$ turbulence at the low-field side (LFS) mid-plane ($\rho\approx 0.95$), measured by the Doppler backscattering systems, shows a steady increase in the root-mean-square (RMS) amplitude (Figs. 6(e) and 6(f)). However, following the onset of the radiative cooling, a dramatic shift occurs: the edge density abruptly saturates (blue in Fig. 6(a)), the electron temperature drops sharply (Fig. 6(b)), and the LFS edge turbulence amplitude also saturates (Fig. 6(d)).
In contrast, line-integrated, low-$k$ density fluctuations at the HFS (Figs. 6(e) and 6(f)) and midplane (Figs. 6(g) and 6(h)) show a different evolution. The fluctuation amplitudes from both the HFS CO2 interferometer and the midplane RIP diagnostic begin to grow substantially after the onset of edge cooling. This divergent behavior—saturation of localized LFS turbulence contrasted with a continued, strong rise in line-integrated HFS and midplane fluctuations—points to a significant poloidal asymmetry in the transport. The turbulence becomes intense in the HFS region where the radiative front develops, linking the enhanced fluctuations and transport to the dynamics of the edge cooling potentially driven by radiative instabilities.
III.3 The Edge Density Limit: Role of Radiative Cooling and Instabilities
III.3.1 Development of Radiation Front at the HFS edge
The evolution of profiles and turbulence suggests the formation of a radiative front and the resulting edge cooling as a potential limiting mechanism for the edge density. To elucidate the role of this radiative instability, the analysis now focuses on the edge region to examine the dynamics of the 2D radiated power distribution.

### v3

# Disentangling Core and Edge Mechanisms of the Density Limit in DIII-D Negative Triangularity Plasmas

@@ Abstract
The density limit is investigated in the DIII-D negative triangularity (NT) plasmas which lack a standard H-mode edge. We find the limit may not be a singular disruptive boundary but a multifaceted density saturation phenomenon governed by distinct core and edge transport mechanisms. Sustained, non-disruptive operation is achieved at densities up to 1.8 times the Greenwald limit ($n_{\mathrm{G}}$) until the termination of auxiliary heating. Systematic power scans reveal distinct power scalings for the core ($n_{e}\propto P_{\mathrm{SOL}}^{0.27\pm 0.03}$) and edge ($n_{e}\propto P_{\mathrm{SOL}}^{0.42\pm 0.04}$) density limits. The edge density saturation is triggered abruptly by the onset of a non-disruptive, high-field side radiative instability that clamps the edge density below $n_{\mathrm{G}}$. In contrast, the core density continues to rise until it saturates, a state characterized by substantially enhanced core turbulence. Core transport evolves from a diffusive to an intermittent, avalanche-like state, as indicated by heavy-tailed probability density functions (kurtosis $\approx 6$), elevated Hurst exponents, and a $1/f$-type power spectrum. These findings suggest that the density limit in the low-confinement regime is determined by a combination of edge radiative instabilities and core turbulent transport. This distinction provides separate targets for control strategies aimed at extending the operational space of future fusion devices.
@@ I Introduction
High plasma density is a fundamental requirement for achieving a self-sustaining fusion reaction, as fusion power scales strongly with density [1]. However, tokamak operation is constrained by an operational density limit [2, 3, 4]. Exceeding this critical density leads to a rapid loss of confinement, often culminating in a plasma termination. Overcoming this long-standing obstacle is therefore a critical step toward realizing economically viable fusion energy.
@@
The conventional metric for the density limit is the empirical Greenwald scaling [2, 3]: $n_{\text{G}}\,(10^{20}\,\mathrm{m^{-3}})=\frac{I_{\text{p}}\,(\text{MA})}{\pi a^{2}\,(\mathrm{m^{2}})}$, where $I_{\text{p}}$ denotes the plasma current and $a$ the minor radius. This scaling is notably independent of heating power. This paradigm is increasingly challenged by observations that the achievable density scales favorably with heating power. Recent theoretical modeling [5, 6] and database analyses [7, 8] suggest that the achievable density limit in the low-confinement (L-mode) regime may increase with heating power, potentially exceeding the conventional Greenwald density. Furthermore, the power-dependent density limits have also been observed in stellarators [9, 10], which lack plasma current, suggesting a more universal physical basis rooted in power balance and transport that transcends specific magnetic configurations. Therefore, further dedicated investigations are needed to develop a physics-based understanding of the fundamental mechanisms governing density limits in magnetically confined plasmas.
@@
The physics underpinning the L-mode density limit involves a complex interplay of several processes. Turbulent transport [11, 12, 13, 14, 15], radiative power losses [16, 7, 4], and magnetohydrodynamic (MHD) instabilities [17, 18] have all been reported as key contributors. These processes do not act in isolation; instead, they are coupled synergistically [3] and ultimately constrain the plasma density. Consequently, developing predictive models of density limits requires disentangling their individual roles and interdependencies.
@@
To elucidate the fundamental physics of the L-mode density limit, it is advantageous to use a plasma configuration that approximates L-mode edge conditions over a wide range of operational parameters. This avoids the complexities introduced by the edge transport barriers (ETBs) and edge localized modes (ELMs), which are characteristic of conventional high-confinement (H-mode) plasmas. The negative triangularity (NT) plasmas [19, 20], characterized by an inverted “D”-shaped cross-section (see Fig. 1), provides an ideal platform for this study. These plasmas exhibit improved core confinement [21, 22, 23, 24] while maintaining L-mode-like edge conditions. While the NT edge is not identical to that of a conventional positive triangularity L-mode, its lack of strong H-mode ETBs and ELMs makes it an excellent proxy for isolating the core and edge transport phenomena that govern density limits without confounding factors.
@@
This paper presents an examination of the physics governing the density limit in NT plasmas on the DIII-D tokamak. We report the achievement of sustained, non-disruptive high-density operation, where the plasma density can surpass the conventional Greenwald density by a factor of 1.8 at substantial input power. A systematic power scan reveals disparate scaling dependencies for the core ($n_{e}\propto P^{0.27\pm 0.03}_{\text{SOL}}$) and the separatrix ($n_{e}\propto P^{0.42\pm 0.04}_{\text{SOL}}$) regions. Detailed experimental measurements reveal that different mechanisms determine density limits in the core and edge plasmas. At the plasma edge, the density is constrained by enhanced turbulence driven by a radiative instability. In contrast, the core density is not directly limited by the radiative cooling but eventually saturates due to the emergence of enhanced, avalanche-like turbulence. These findings challenge the traditional paradigm of a single disruptive limit, advancing a more nuanced picture where distinct processes govern density saturation in different plasma regions.
@@ II Experimental Arrangement
The experiment was conducted on the DIII-D tokamak [25], with auxiliary heating power provided by neutral beam injection (NBI) and varied between discharges. All discharges employed a lower single-null configuration with an open divertor geometry, and the ion $\mathbf{B}\times\nabla\mathbf{B}$ drift was directed away from the primary divertor, as illustrated in Fig. 1. The plasma shape was characterized by negative triangularity at both the top and bottom ($\delta_{\text{top}}=-0.3$ and $\delta_{\text{bot}}=-0.62$), moderate elongation ($\kappa\approx 1.3$), and an edge safety factor of $q_{95}=4.1$. This paper focuses specifically on discharges with toroidal field $B_{t}=2.0$ T and plasma current $I_{p}=0.6$ MA. Additional experiments exploring different plasma current and toroidal field combinations are reported separately [26]. In each discharge, plasma density was gradually increased via gas puffing from the top of the mainchamber. Consistent with the behavior of NT plasmas, no edge localized modes were observed in any discharge.
@@
A comprehensive suite of diagnostics monitored the evolving plasma state. Figure 1 shows the locations of the key diagnostics. Electron density and temperature in the main chamber were measured using the Thomson scattering (TS) system [27]. Charge exchange recombination (CER) spectroscopy [28] measured the temperature, density, and toroidal and poloidal rotation velocities of carbon ions. These impurity ion profiles were used to compute radial electric field profiles ($E_{r}$) using the carbon ion force balance equation. Bolometer arrays [29] provided measurements of the radiated power distribution and total radiated power.
Several diagnostics were used to probe density fluctuations. A multichannel $\text{CO}_{2}$ interferometer [30] measured the line-averaged density and its low-$k$ fluctuations along one horizontal midplane chord and three vertical chords at major radii of $R=1.48\,\mathrm{m}$ (high-field side), $R=1.94\,\text{m}$ (core region), and $R=2.10\,\text{m}$ (low-field side). The radial interferometer polarimeter (RIP) [31] also provided line-integrated measurements of electron density fluctuations at the mid-plane. A 2D array of high-sensitivity beam emission spectroscopy (BES) [32] measured ion-scale ($k_{\theta}<3\,\mathrm{cm^{-1}}$) density fluctuations at the edge. Two 8-channel Doppler back-scattering (DBS) systems [33, 34] measured turbulent flows and low-to-intermediate-$k$ ($2<k_{\theta}<10\,\mathrm{cm^{-1}}$) density fluctuations. The radial locations and corresponding wavenumbers of the DBS measurements were calculated using the 3D ray-tracing code GENRAY [35].
@@ III.1.1 Density Saturation Beyond the Greenwald Limit
A key result of this study is the achievement of sustained L-mode operation at densities that surpass the conventional Greenwald density limit. This is shown in Fig. 2, which presents the time evolution of key parameters in two representative discharges with different levels of auxiliary heating from NBI. Across the experiment, the NBI power was varied from 3 MW up to 13 MW. As shown in Figs. 2(b) and 2(c), both core and separatrix densities increase with input power. The core density at $\rho\approx 0.7$, which is similar to the line-averaged density, consistently exceeds the Greenwald density ($n_{\text{G}}$). Notably, rather than disrupting upon reaching their maximum density, the discharges with medium to high NBI power ($P_{\mathrm{NBI}}>4~{}\mathrm{MW}$) transitioned into a phase of density saturation that persisted until the termination of auxiliary heating.
The radiated power in the divertor [Fig. 2(d)] and the main chamber [Fig. 2(e)] also increase with density, indicating enhanced radiative losses. The spikes in the divertor radiation power [Fig. 2(d)] mark the development of a radiation front that propagates from the divertor towards the high-field side (HFS).
The line-averaged density, inferred from Thomson scattering profiles, reached $1.8\times n_{\text{G}}$ in the high-power discharge. For comparison, the CO2 interferometer measured a line-averaged density of $2.0\times n_{\text{G}}$ [20, 26]. This discrepancy arises because the interferometer’s measurement includes the substantial scrape-off layer density present during high-density operation. Therefore, Thomson scattering measurements are used for the analysis in this study.
@@
III.1.2 Disparate Power Scaling for Core and Edge Density Limit
Having established that the density limit in these NT plasmas is a saturation phenomenon, we investigate its dependence on heating power. The analysis correlates the maximum achieved density with the power flowing into the scrape-off layer ($P_{\mathrm{SOL}}$), defined as $P_{\mathrm{SOL}}=P_{\mathrm{in}}-P_{\mathrm{rad,core}}-\mathrm{d}W/\mathrm{d}t$, accounting for total input power, core radiated power, and changes in stored energy. This parameter represents the effective supplied heating. All values are averaged over a 100 ms window at the point of maximum density.
A systematic power scan reveals that the core and edge densities respond differently to $P_{\mathrm{SOL}}$, as shown in Fig. 3. The maximum core density (at $\rho=0.7$) exhibits a weak but clear power dependence, scaling as $n_{e,\text{core}}\propto P_{\mathrm{SOL}}^{0.27\pm 0.03}$. This result is consistent with L-mode density limit scalings reported on other tokamaks, such as ASDEX Upgrade [8]. In contrast, the separatrix density displays a stronger relationship, scaling as $n_{e,\text{sep}}\propto P_{\mathrm{SOL}}^{0.42\pm 0.04}$. This relatively strong power dependence is reminiscent of the density limit observed in stellarators, where power balance is known to be the dominant constraint [9, 10].
These disparate power scalings constitute a central finding of this work; they indicate that the density saturation is not governed by a single, global mechanism. Instead, it suggests that distinct physical processes, with different sensitivities to heating power, are responsible for limiting the density in the core and edge regions. This hypothesis—that core saturation is dominated by turbulent transport while edge saturation is governed by radiative instabilities—guides the investigative framework for the remainder of this paper.
@@ III.1.2 Disparate Power Scaling for Core and Edge Density Limit
III.2 Profile Evolution and Confinement Degradation Towards Density Limit
III.2.1 Reduced Temperature and Rotational Shear at High Density
As the plasma approaches the density limit, its equilibrium profiles evolve in a manner that signals a clear degradation of thermal confinement. This evolution, shown for a representative high-power discharge in Fig. 4, provides important insight into the underlying transport dynamics.
@@ III.2.1 Reduced Temperature and Rotational Shear at High Density
The electron density profile [Fig. 4(a)] rises across the radius but remains notably flat in the core, indicating that strong particle transport persists even at high densities. The degraded confinement is accompanied by a loss of thermal energy, as both the electron and ion temperature profiles collapse [Figs. 4(b), 4(c)]. A noticeable feature is the pronounced flattening of the edge $T_{e}$ profile, a classic signature of intensified radiative cooling at the plasma boundary.
Another critical change occurs in the plasma rotation. The toroidal rotation profile [Fig. 4(d)], initially peaked in the core, progressively collapses with increasing density. This loss of core rotation directly dismantles the dominant component of the radial electric field ($E_{r}$), a key factor for turbulence suppression. As shown in Fig. 5, the core $E_{r}$ shear, initially strong, diminishes and ultimately vanishes as the density limit is approached.
@@
This cascade of events, from intensified edge cooling to the collapse of core rotation and $E_{r}$ shear, indicates that the plasma’s ability to suppress turbulence deteriorates with rising density. The loss of this stabilizing flow shear enables enhanced turbulent transport that governs the core density limit, as subsequent sections will demonstrate.
@@ III.2.2 Divergent Evolution of Turbulence in Different Regions
The degradation of thermal confinement described above coincides with a dramatic divergence in turbulent activity across the plasma, as illustrated in Fig. 6 for a high-power discharge. A critical transition occurs at approximately 3700 ms (vertical dashed line), marking the onset of strong edge cooling. Before this event, the turbulence evolution is relatively uniform; afterward, it diverges sharply between different spatial regions.
@@
Initially, both core and edge densities rise together [Fig. 6(a)], accompanied by a steady increase in the amplitude of local, intermediate-$k$ turbulence at the low-field side (LFS) midplane ($\rho\approx 0.95$), as measured by Doppler backscattering [Fig. 6(d)]. This period of evolution ends abruptly at the 3700 ms transition. Immediately following this transition, the LFS turbulence saturates. This saturation coincides with the clamping of the edge density and the sharp drop in local electron temperature [Fig. 6(b)].
In contrast, turbulence in other regions intensifies. Line-integrated, low-$k$ density fluctuations on the high-field side (HFS), measured by the CO2 interferometer, grows substantially [Figs. 6(e), 6(f)]. A similar growth is observed in the line-integrated midplane fluctuations measured by the RIP diagnostic [Figs. 6(g), 6(h)].
This divergent evolution—saturation of localized turbulence on the LFS versus continued, strong growth of HFS and core fluctuations—reveals a profound poloidal asymmetry in the transport dynamics. This asymmetry strongly suggests that the density limit is not a globally uniform process but is instead triggered by a localized instability on the high-field side—likely the radiative phenomenon that we investigate next.
III.3 The Edge Density Limit: A Radiatively Driven Phenomenon
III.3.1 Visualizing the HFS Radiation Front
The poloidally asymmetric turbulence described previously points to a localized radiative instability as the trigger for the edge density limit. Here, we examine the 2D radiated power distribution, reconstructed from bolometer measurements, to visualize the formation and evolution of this instability. Figure 7 presents a time-sequence of this evolution. While geometrically similar to a disruptive MARFE [16, 36], the feature observed here is a quasi-stable, non-disruptive radiation front that is directly linked to the observed density saturation.